OpenStudy (anonymous):

The perimeter of a retangle is 40m. The length is 8m more than twice the width. find the demensions. what is the lenght and width please?

OpenStudy (anonymous):

you know that the length is 8m more than twice the width:\[L = 8 + 2x\] and \[W = x\] The perimeter of a rectangle is:\[P = 2L + 2W\] hence,\[40 = 2(8+2x) + 2(x)\]\[40 = 16 + 4x + 2x\]\[24 = 6x\]\[x = 4\] Therefore,\[L = 8+2(4) = 16m\] and \[W=4m\]

OpenStudy (anonymous):

oK sO I need to to put 16 for the lenght and 4 width